# Top-of-atmosphere radiation fields over the last millennium reconstructed from proxies

### v1

@@ 4.5 Volcanic eruptions
We composite pre-industrial tropical eruptions larger than Pinatubo, i.e., eruptions with a volcanic stratospheric sulfur injection (VSSI) of at least 7.5 Tg S and a latitude below 30°. If a second eruption with a VSSI >> 1 Tg S occurs within the composite window, we remove all data following that second eruption to capture the undisturbed response. Eruption dates, latitude, and VSSI are taken from the eVolv2k dataset (Toohey and Sigl, 2017), which is derived from ice core records.
There is a clear global cooling signal, albeit with large scatter, peaking one to two years after the eruption at 0.15-0.15 °C (NH: 0.22-0.22 °C, SH: 0.09-0.09 °C) and fully recovering after 10–15 years (Fig. 10). The stronger NH cooling, particularly in summer (Fig. S12a), is possibly due to proxy distribution and seasonality, but model simulations also suggests a hemispherically asymmetric response (Pauling et al., 2021). The OHC300 is reduced by 22 ZJ after five years, although most heat is lost in the first two years after the eruption. This heat loss is driven by enhanced RSR (up to +0.7+0.7 W m-2), counteracted by a small reduction in OLR (up to 0.3-0.3 W m-2). This leads to a negative EEI of up to 0.4-0.4 W m-2, which approximately aligns with the rate of OHC loss. Most OHC is lost over the north Atlantic and the equatorial west Pacific (Fig. S13). EEI becomes positive after 5–10 years, coinciding with the recovery of OHC. Time scales for sea ice expansion are similar to those of surface temperature, and there is no discernible impact on ENSO (Fig. S12), consistent with the analysis of Zhu et al. (2022). Compared to coupled-model simulations, the volcanic response in all fields is damped and lagged (Fig. S14).
Figure 10: Composite analysis of global-mean seasonal anomalies for tropical volcanic eruptions larger than Pinatubo (1991) over 850–1850 (n = 19). Eruptions have a VSSI of at least 7.5 Tg S, a latitude below 30°, and are marked in orange in Fig. 5a. Anomalies are computed relative to the 5 years prior to each eruption. The blue and green lines in (b) correspond to eruptions causing large OHC losses. The orange line in (d) is the 3-point moving average of the time derivative of OHC300, rescaled by Earth’s ocean fraction (71%) to make it comparable to EEI.
Multiple large eruptions spaced a few years to a decade apart can initiate periods of prolonged OHC loss (Fig. S15, and colored line in Fig. 10b; Gupta and Marshall, 2018). For example, the 1693 and 1695 eruptions caused a 10-year period of OHC loss. After a loss of 55 ZJ, the recovery took 40 years due to five additional, smaller eruptions.
@@ 4.6 Consistency between OHC and EEI
Approximately 90% of the heat absorbed due to a non-zero EEI is stored in the ocean (von Schuckmann et al., 2023). Recently, 42% of the accumulated OHC has been stored in the upper 300 m (Pan et al., 2025). Therefore, a strong correlation between the EEI and the rate of change of the upper-ocean heat content (dOHC300/dt) is expected, particularly at seasonal to interannual timescales. This can provide insight into the connection between Earth’s energy budget at the TOA and at the surface, but also serves as a consistency check for the DA procedure. By Eq. (1), a difference between dOHC/dt $\approx C(dT/dt)$ and the EEI may be due to heat uptake $\gamma T$ by the deeper ocean, or components of the climate system other than the ocean. We multiply dOHC/dt by Earth’s ocean fraction (71%) to obtain the equivalent EEI over the total surface area. The derivative is calculated using a central difference as in  (Trenberth et al., 2014).
Reconstructed EEI and dOHC300/dt are correlated annually at 0.69 (Fig. S16), and the volcanic response is similar (Fig. 10d). At seasonal timescales, the correlation is 0.60, but dOHC/dt leads EEI by one season. With increasing timescale, the two quantities decorrelate. Interpreting the difference as deep ocean heat uptake, the deep ocean is releasing heat over most of the last millennium.
5 Discussion and Conclusions
We presented a reconstruction of Earth’s energy budget over the last millennium, comprising both the TOA radiation and upper-ocean heat content, in addition to radiatively relevant surface variables. In contrast to previous reconstructions, ours is spatially and seasonally resolved at the surface and at the TOA, separates the SW and LW components, and is constrained by proxies. The reconstruction can be used to investigate seasonal to decadal variability, millennial trends, and episodic events like volcanic eruptions.
@@ 5.1 Reconstruction skill and limitations
The reconstruction presented here is the first proxy-constrained dataset of TOA radiation over the last millennium. We use the fact that radiation covaries with surface temperature, to which the proxies are sensitive. However, the proxy information is temporally and spatially sparse, and the link to TOA radiation is mediated by complex and noisy physical processes. Therefore, we rely on a pseudoproxy experiment, instrumental validation, and internal consistency to establish TOA radiation reconstruction skill. As in previous reconstructions, surface variables agree well with instrumental products and other reconstructions (Figs. S3–S5), and withheld proxies for independent validation over the whole last millennium have significant positive correlations (Fig. S6).
Of the three TOA radiation fields, OLR has the highest correlations in the PPE and with the ERBE satellite record for the real-proxy reconstruction (Figs. 2 and 3). This is not surprising since OLR covaries strongly with TAS through the Planck response, especially in association with ENSO and other modes of internal variability. RSR has slightly lower correlations because it is generally noisier at seasonal timescales. EEI skill may be lower because it is the small difference of two opposing quantities of similar magnitude. While there is skill at seasonal and annual timescales in all fields, it is higher at interannual to decadal timescales.
@@
Consistency between the TOA radiation and OHC, even though not explicitly enforced, is further evidence of reconstruction skill (Figs. S16 and 10d). EEI and dOHC300/dt should correspond since much of the energy imbalance is absorbed into the ocean (von Schuckmann et al., 2023; Palmer and McNeall, 2014). Indeed, their seasonal and annual correlation is high, with a correlation peak at a timescale of 4 years. Up to timescales of 20 years, the correlation is above 0.60, with a trend toward decorrelation for longer timescales. We speculate that the multidecadal correlation might be higher if OHC were to include the deeper ocean. At seasonal timescales, the EEI lags dOHC/dt by one season.
One caveat of our method is that the partitioning of EEI into RSR and OLR only works well over the last millennium, and not over the historical period. This is because the temperature–radiation relationship in the LIM is learned from last-millennium simulations, which are dominated by volcanic forcing rather than greenhouse gases and tropospheric aerosols. In the training data, a cooling (warming) is thus usually associated with a negative (positive) SW forcing, which in response reduces (increases) the OLR. This is in contrast to anthropogenic greenhouse gas forcing, for which warming is associated with reduced OLR, opposite of what the LIM has learned. Interannual variability over the historical period does not suffer from this issue, however, as evident from the satellite comparison (Fig. 3).
@@
We conclude that our reconstruction is generally skillful, despite the spatiotemporal sparseness of the proxy network and the complex relationship between temperature and radiation. Most reliable are global means at interannual to decadal timescales, although even seasonal fields have considerable skill. The spatial reconstruction fidelity is highest in the tropical and subtropical Pacific, but non-negligible in most regions. Over the historical period, the partitioning at decadal timescales is inaccurate, but the total EEI and non-radiation fields do not lose skill.
@@ 5.2 Earth’s energy budget over the last millennium
Reconstructions and models agree that there was a global-mean cooling trend over the last millennium, although the patterns of cooling may not be globally synchronous (PAGES 2k Consortium, 2019; Neukom et al., 2019). This cooling was driven primarily by volcanic forcing, with a potential minor role of greenhouse gas and solar forcing (Schurer et al., 2014; Büntgen et al., 2020; Wanner et al., 2022; Esper et al., 2012), then amplified by positive feedbacks (Atwood et al., 2016). However, the climate system is stable due to the strongly negative Planck feedback, which responds to cooling with reduced OLR. Indeed, the OLR decreased by 0.42 W m-2 kyr-1, while the RSR only increased by 0.29 W m-2 kyr-1, which explains the increasing EEI trend of 0.16 $\pm$ 0.26 W m-2 kyr-1, although it is not significantly different from zero. To sustain the cooling trend, the EEI needs to remain negative on average despite the increasing trend, leading to a slowdown of cooling over centuries. In our reconstruction, this implies an 1850–1900 EEI mean of at most 0.15 $\pm$ 0.25 W m-2 (cf. Fig. S9b). If we let the relatively constant SAT and OHC over 1600–1900 imply an average EEI of zero over this period, the 1850–1900 EEI mean would be 0.07 $\pm$ 0.21 W m-2, similar to the estimate of 0.2 $\pm$ 0.2 W m-2 from Forster et al. (2021). In comparison, the EEI inferred from the OHC reconstruction of Gebbie and Huybers (2019) is 0.24 W m-2 over this period. The energy imbalance over any 20-yr period after 1980, and possibly after 1970, is unprecedented in the context of the last millennium, similar to temperatures (PAGES 2k Consortium, 2019). This provides further evidence that the global warming that started in the 20th century is anthropogenic rather than natural (Masson-Delmotte et al., 2021).
While both our reconstruction and that of Gebbie and Huybers (2019, described further in , ) show sustained energy loss before around 1600, theirs starts with a higher energy content and cools more into the LIA (Fig. S9). The discrepancy may be explained by the much colder pre-industrial SSTs prescribed by Gebbie and Huybers (2019) than we reconstruct (Fig. S8). Their reconstruction then agrees well with the reconstruction from Wu et al. (2025) over the historical period since both are instrumental-based, while our EEI is likely too low in the early 1900s, even though it matches in the later decades. Our reconstructions agree that the energy inventory was higher during the MCA than in 2000, and that there was significant multidecadal variability with intermittent periods of energy gain. However, both reconstructions come with large uncertainties that preclude definite conclusions about the last millennium energy budget.
@@
By comparing the spatial structure of the radiation trends (Figs. 9 and S11), we can infer the physical processes involved. The OLR trend is negative almost everywhere, presumably due to the Planck feedback. Superimposed is a La Niña-like response (cf. Fig. 1), with a westward shift of the Indo–Pacific convection region toward the Maritime Continent, causing negative OLR trends in regions with enhanced deep convection (Hartmann, 2016). The increased EEI south of the convective region points to a southward shift of the Intertropical Convergence Zone accompanying the westward shift, likely because the SH cools less and thus more heat is transported northward (e.g., Donohoe et al., 2013). In addition to these shifts, the RSR and EEI have trends in the subtropical east Pacific and Atlantic, likely associated with the subtropical stratus decks. Northeast Pacific low cloud cover appears enhanced and thus more reflective, while southeast Pacific and Atlantic low clouds appear somewhat suppressed. There also is a clear-sky RSR contribution from enhanced Antarctic sea ice (Fig. S11a).
5.3 Seasonality due to orbital precession
The last-millennium trends of all fields differ consistently by season. The seasonal cycle in the global-mean cooling trend ranges from 0.27-0.27 K kyr-1 in MAM to 0.32-0.32 K kyr-1 in SON (Fig. 8). The timing is latitude-dependent since insolation trends are most extreme in MMAJJA in the NH and SONDJF in the SH (Fig. 7c,d; cf. Fig. 2d in Lücke et al., 2021). Sea ice area trends are synchronous with these high-latitude insolation trends, with lower insolation coinciding with more sea ice. These seasonal trends are associated with a reduced (enhanced) amplitude of the seasonal sea ice cycle in the Arctic (Antarctic). The same pattern is also evident in the radiation fields (Fig. 8d–f).
The insolation trends are caused by axial precession (Lücke et al., 2021). While the annual-mean insolation is approximately constant over the last millennium, there are latitude-dependent seasonal trends in insolation due to shifting of the solstices relative to the apsides of Earth’s orbit around the sun. MAM thus tends to occur closer to perihelion, attenuating the cooling, while SON tends to occur closer to aphelion, amplifying the cooling (Fig. 8a). This has previously led to a wrong interpretation of trends in annual proxy reconstructions (Lücke et al., 2021), while our seasonal DA approach explicitly accounts for such seasonality.

### v2

@@ 4.5 Volcanic eruptions
We composite pre-industrial tropical eruptions with a volcanic stratospheric sulfur injection (VSSI) of at least $9\,\mathrm{Tg\ S}$ (similar to Mt. Pinatubo in 1991) and a latitude equatorward of 30°. For volcanic events within 10 years of each other, we only use the second eruption. Eruption dates, latitude, and VSSI are taken from the eVolv2k version 4 database (Toohey and Sigl, 2017; Sigl and Toohey, 2024), which is derived from ice core records.
Figure 10: Composite analysis (also known as superposed epoch analysis) of global-mean seasonal anomalies for large tropical volcanic eruptions over 850–1850 (n = 14). Eruptions have a VSSI of at least $9\,\mathrm{Tg\ S}$, a latitude equatorward of 30°, and are marked in orange in Fig. 5a. Anomalies are computed relative to the 5 years prior to each eruption, and only the second eruption is used for volcanic double events. The green line in (b) corresponds to the 1695 eruption. The orange line in (d) is the composite of the time derivative of OHC300, rescaled by Earth’s ocean fraction (71%) to obtain the equivalent EEI and smoothed with an annual running mean. This is expected to differ from the time derivative of the composite of OHC300 in (b) since the pre-eruption climatology is different.
There is a clear global cooling signal, albeit with large scatter across individual eruptions, peaking one to two years after the eruption at $0.17\,^{\circ}\mathrm{C}$ (NH: $0.24\,^{\circ}\mathrm{C}$, SH: $0.12\,^{\circ}\mathrm{C}$) on average and fully recovering after 10–15 years (Fig. 10). The stronger NH cooling, particularly in summer (Fig. S15a), is possibly due to proxy distribution and seasonality, but model simulations also suggest a hemispherically asymmetric response (Pauling et al., 2021). Cooling is within the spread of the proxy reconstruction and model composite by He et al. (2026). OHC300 is reduced by around $25\,\mathrm{ZJ}$ on average, most rapidly during the first two years after the eruption. This heat loss is coincident with enhanced RSR (up to $0.8\,\mathrm{W\,m^{-2}}$), a small reduction in OLR (up to $0.6\,\mathrm{W\,m^{-2}}$) and negative EEI of up to $0.2\,\mathrm{W\,m^{-2}}$. EEI becomes positive after 3–5 years, and the OHC recovers over the course of a decade. Sea-ice expansion has a temporal response similar to surface temperature, and there is no discernible impact on ENSO (Fig. S15), consistent with other reconstructions and models (He et al., 2026; Zhu et al., 2022; Dee et al., 2020; Dee and Steiger, 2022). Compared to coupled-model simulations, the volcanic response in all fields is muted and time-lagged (Fig. S16). Relative to the reconstruction, model-simulated cooling is three times larger, and tTOA radiation anomalies are 10–20 times larger.
Multiple large eruptions spaced a few years apart can initiate periods of prolonged OHC loss (Fig. S17; Gupta and Marshall, 2018). For example, the 1693 and 1695 eruptions caused a decade of OHC loss, peaking at $70\,\mathrm{ZJ}$ (green line in Fig. 10b). The recovery from these two eruptions took 30 years due to five additional, smaller eruptions.
@@ 4.6 Consistency between OHC and EEI
Approximately 90% of the heat absorbed due to a non-zero EEI is stored in the ocean (von Schuckmann et al., 2023). Recently, 42% of the accumulated OHC has been stored in the upper 300 m (Pan et al., 2025). Therefore, a strong correlation between the EEI and the rate of change of the upper-ocean heat content (dOHC300/dt) is expected, particularly at seasonal to annual timescales, as described by Eq. (1). This can provide insight into the connection between Earth’s energy budget at the TOA and at the surface, but also serves as a consistency check for the DA procedure. We multiply dOHC/dt by Earth’s ocean fraction (71%) to obtain the equivalent EEI over the total surface area. The derivative is calculated using a central time difference as in Trenberth et al. (2014).
Reconstructed EEI and dOHC300/dt are correlated annually at $r=0.64$ ($p\leq 0.01$; Fig. S18). The volcanic response is similar in shape and timescale, but EEI is weaker than dOHC300/dt (Fig. 10d). At seasonal timescales, the correlation is $r=0.50$ ($p\leq 0.01$), but dOHC/dt leads EEI by about one season. With increasing timescale beyond four years, the two quantities decorrelate (Fig. S18e).
5 Discussion
Our reconstruction of Earth’s energy budget over the last millennium represents a new application of paleoclimate DA. After establishing the reconstruction skill and its limitations, this new dataset allows us to consider several aspects of natural energy budget variability, constrained by proxy information and model dynamics.
@@ 5.1 Reconstruction skill and limitations
The reconstruction presented here is the first proxy-constrained dataset of TOA radiation over the last millennium. We use the fact that radiation covaries with surface temperature, to which the proxies are sensitive. However, the proxy information is temporally and spatially sparse, and the link to TOA radiation is mediated by complex and noisy physical processes. Therefore, we rely on pseudoproxy experiments, instrumental validation, and internal consistency to establish TOA radiation reconstruction skill. As in previous reconstructions, surface variables generally agree well with instrumental products and other reconstructions (Figs. S7–S9 and S11), and withheld proxies for independent validation over the whole last millennium have a significantly positive network-mean correlation (Fig. S6). However, Antarctic SIC is likely unreliable, and NH extratropical land summer temperatures are too cold.
The energy imbalance at interannual timescales can be reliably reconstructed. This follows from the PPEs (Figs. 2 and S4) and the satellite comparison (Fig. 3). Interannual variability is dominated by the radiative response to unforced surface temperature variability (Trenberth et al., 2015), to which our proxies are sensitive. Indeed, our PPEs showed that the radiative response can be reconstructed accurately from proxies (Fig. S4 and Table S1), explaining the high interannual EEI skill. The magnitude of volcanic forcing, another important but sporadic source of interannual variability, however, is greatly underestimated. This is the case both in PPEs (Figs. 2 and S4) and in the real reconstruction (Fig. 10). Nonetheless, a volcanic signal is present and the relationship to other variables is physically consistent. In particular, our DA system is able to capture the complex lead/lag relationship of forcing and response with temperature (Fig. S4d; Proistosescu et al., 2018).
The constituent RSR and OLR of the energy imbalance have good global-mean skill in the PPE (Fig. 2). In the satellite comparison, correlations are not significant, although this is very likely a result of the short overlap period (Fig. 3). Spatial OLR skill is higher than spatial RSR skill. This is not surprising since OLR covaries strongly with SAT through the Planck response and is spatially coherent, especially in association with ENSO and other modes of internal variability. RSR has slightly lower correlations because it is generally noisier at seasonal timescales due to extratropical storms and the Madden–Julian oscillation (e.g., Trenberth et al., 2015). In the PPE, the EEI skill is lower than RSR and OLR, possibly because the EEI is the small residual of two large quantities.
@@
Consistency between the TOA radiation and OHC, even though not explicitly enforced by the DA system, is further evidence of reconstruction skill (Figs. S18 and 10d). EEI and dOHC300/dt should correspond since much of the energy imbalance is absorbed into the ocean (von Schuckmann et al., 2023; Palmer and McNeall, 2014; Allison et al., 2020). Indeed, their seasonal and annual correlation are moderately high ($r=0.5$ to $0.6$), with a correlation peak at a timescale of 4 years ($r=0.7$). The correlations between EEI and dOHC/dt are similar for instrumental estimates, and would likely be higher if atmospheric heat storage were considered (Hakuba et al., 2024; Mayer et al., 2026). Up to timescales of 20 years, the correlation is above $0.6$, with a trend toward decorrelation for longer timescales. We speculate that the multidecadal correlation might be higher if OHC were to include the deeper ocean. At seasonal timescales, the EEI lags dOHC/dt by one season (Fig. S18d). This may be a result of the more direct covariance of OHC with proxies, while EEI mostly follows from LIM dynamics, or an indication that much of the interannual variability in EEI results from internal fluctuations in OHC rather than external forcing (cf. Proistosescu et al., 2018).
One caveat of our method is that we can only reconstruct the radiative response and natural forcing, but not the anthropogenic forcing dominating the historical period. This is because the temperature–radiation relationship in the LIM is learned from last-millennium simulations, which do not include a strong greenhouse gas forcing signal since they end in 1850. A nonlinear reconstruction framework may be required to capture relationships that change over time and with forcing type. However, if we add an estimate of the anthropogenic forcing to our reconstructed EEI, the sum agrees well in magnitude and phasing with an OHC-based estimate of the total EEI (Fig. 6). Only after 1960 do they diverge as a result of an underestimated radiative response in our reconstruction.
The reconstruction skill diminishes as proxies become more sparse toward the early part of the last millennium. Additionally, there may be a drop in skill before 1100, when no coral proxies are available to constrain the tropical Pacific. This drop is evident in PPEs but not in the withheld-proxy correlations. Further, our reconstruction has a last-millennium warming trend over North America and parts of continental Asia, particularly in summer (Fig. S14a). This trend may be spurious and would explain the negative correlations for some TRW proxies before 1600 (Fig. S6d). During the development of our DA system, we also experimented with other PSM calibration datasets and periods. While the presented version agrees best with instrumental temperatures, other versions tended to have more temperature variance, resulting in a colder LIA and a steeper last-millennium cooling trend. We suggest that PSM calibration uncertainties should be considered in future DA reconstructions, for example, by sampling calibration datasets and periods across Monte Carlo iterations, and by making use of recent instrumental temperature ensembles (e.g., Chan et al., 2024; Lenssen et al., 2024).
@@
We conclude that the reconstruction is skillful for surface fields, the radiative response associated with internal variability, and the timing of volcanic eruptions, despite the spatiotemporal sparseness of the proxy network and the complex relationship between temperature and radiation. However, the magnitude of volcanic forcing is greatly underestimated, and anthropogenic forcing is not represented. Most reliable are global means at interannual to decadal timescales after 1100, although even seasonal fields have considerable skill. The reconstruction fidelity is highest in the tropics and subtropics.
@@ 5.2 Earth’s energy budget over the last millennium
Reconstructions and CMIP5 models agree that there was a global-mean cooling trend over the last millennium, although the patterns of cooling may not have been globally synchronous (PAGES 2k Consortium, 2019; Atwood et al., 2016; Neukom et al., 2019). This cooling was driven primarily by volcanic forcing, with a potential minor role of greenhouse gas and solar forcing (Schurer et al., 2014; Büntgen et al., 2020; Wanner et al., 2022; Esper et al., 2012), then amplified by positive feedbacks (Atwood et al., 2016). Since the climate system is overall stable due to the strongly negative Planck feedback, it responds to cooling with reduced OLR. Indeed, the OLR decreased by $0.35\,\mathrm{W\,m^{-2}\,kyr^{-1}}$, while the RSR only increased by $0.16\,\mathrm{W\,m^{-2}\,kyr^{-1}}$, which explains the increasing EEI trend of $0.18\pm 0.21\,\mathrm{W\,m^{-2}\,kyr^{-1}}$, although it is not significantly different from zero (Fig. 8). To sustain the cooling trend, the EEI needs to remain negative on average despite the increasing trend, leading to a slowdown of cooling over centuries. Indeed, both our reconstruction and the OHC reconstruction of Gebbie and Huybers (2019) have a 1100–1850 mean EEI of $0.1\,\mathrm{W\,m^{-2}}$, and a total integrated energy loss over 850–1850 of $1600$ to $1800\,\mathrm{ZJ}$ (Fig. S12). For comparison, Earth has gained about $600\,\mathrm{ZJ}$ over the much shorter 1880–2020 period (Wu et al., 2025).
While the anthropogenic origin of the modern global warming is well-established (Masson-Delmotte et al., 2021), it is merely a “symptom” of a positive energy imbalance (von Schuckmann et al., 2016). By extending the record of energy budget variability to 850, our reconstruction reveals that the current rate of energy gain, which is the fundamental driver of global climate change (Mauritsen et al., 2025; von Schuckmann et al., 2016; Trenberth et al., 2014), is unprecedented in the last millennium. Specifically, the energy imbalance over any 20-yr period after 1990, and possibly already after 1915, exceeds any 20-yr average over 850–1850. Moreover, the 2001–2025 energy imbalance trend is unprecedented among all 25-yr EEI trends in our reconstruction, and is in the 97th percentile of last-millennium model variability (Section 44.2).
@@
By comparing the spatial structure of the radiation trends (Fig. S14d–f), we can speculate on the physical processes involved. The OLR trend is negative almost everywhere, presumably due to the Planck feedback. Superimposed is an El Niño-like response, with an eastward shift of the Indo–Pacific convection region toward the Central Pacific (Hartmann, 2016), causing negative OLR trends in regions with enhanced deep convection and positive OLR trend over the West Pacific Warm Pool. The shift is likely driven by west Pacific cooling, which moves the ascending branch of the Walker circulation to the east (Fig. S14b). In addition to this shift, the EEI has trends in the subtropical east Pacific, likely associated with the subtropical stratus decks. There is also a clear-sky RSR contribution from anomalous Antarctic sea-ice area (Fig. S13a).
Gebbie and Huybers (2019) provide the only other proxy-based reconstruction of the last-millennium energy budget. They propagate SSTs into the ocean interior using a constant circulation model. Annual-mean proxy SSTs from the Ocean2k database (McGregor et al., 2015) are used before 1870, and instrumental winter SSTs after. The SSTs are prescribed as global-mean values rather than spatially resolved fields, but the reconstruction is constrained by basin-wide averages of ship-based vertical temperature profiles in the 1870s. Their SSTs are systematically colder over the last millennium and start to warm about 150 years before most instrumental datasets and our reconstructed SSTs (in 1750 rather than 1900; Fig. S11d), although recent corrections to instrumental data may support a colder 19th century (Chan et al., 2024, 2025). The cooler SSTs also imply a lower OHC (Fig. 5d), but the pre-1850 energy imbalance is not affected by such a mean offset. Our EEI reconstruction is not significantly correlated with theirs, but the mean EEI is similar (Fig. S12). The lack of correlation and the offset in pre-industrial SSTs may be due to differences in proxy network (e.g., the lack of tropical Pacific proxies in Ocean2k), the prescription of global-mean rather than spatially resolved SSTs, the method of proxy-to-temperature conversion, and the blending of winter and annual SSTs. Ocean2k, on which Gebbie and Huybers (2019) is based, may not provide a suitable dataset for annual values in temperature units, rather than the 200-yr-binned standard deviations of proxy units presented in McGregor et al. (2015). However, by resolving the deep ocean, Gebbie and Huybers (2019) may be able to better constrain the energy imbalance on long timescales.
5.3 Seasonality in last-millennium trends
The last-millennium trends of all fields differ consistently by season. The seasonal cycle in the global-mean cooling trend ranges from $0.16\,\mathrm{K\,kyr^{-1}}$ in MAM and JJA to $0.19\,\mathrm{K\,kyr^{-1}}$ in SON and DJF (Fig. 8). High-latitude temperature and insolation trends have a lag correlation, with higher insolation corresponding with warmer temperatures (Fig. 9). The lag of around one season may be a reconstruction artifact, but it could also be a real property of the high-latitude climate, e.g., due to thermal inertia. In the Arctic seasonal cycle, the annual maximum temperature is nearly synchronous with maximum insolation, but the annual minimum temperature lags insolation by around two months (Donohoe et al., 2020).
While the annual-mean insolation is approximately constant over the last millennium, there are latitude-dependent seasonal trends in insolation due to axial precession (cf. Fig. 2d in Lücke et al., 2021). Proxies will reflect these seasonal trends based on their location and seasonality; seasonal DA explicitly accounting for this fact is therefore is essential to separating seasonal and annual-mean trends.